What is the pH of a 0.402 M aqueous solution of NaCH3COO? Ka (CH3COOH) = 1.8x10-5
Answer #1

                    CH3COONa ----------> CH3COO- + Na+

                0.402M                             0.402M

             CH3COO-   + H2O --------> CH3COOH + OH-

    I         0.402M                                 0                0

C          -x                                         +x               +x

   E        0.402-x                                 +x                +x

           kb   = Kw/Ka

                  = 1*10-14/1.8*10-5   = 5.6*10-10

     Kb   = [CH3COOH][OH-]/[CH3COOH]

   5.6*10-10   = x*x/0.402-x

5.6*10-10 *(0.402-x) = x2

     x = 1.5*10-5

   [OH-] = x=1.5*10-5 M

POH = -log[OH-]

             = -log1.5*10-5

            = 4.8239

PH   = 14-POH

         = 14-4.8239   = 9.1761